# The calculus of Duistermaat’s triple index

### v2

@@ 1. Introduction
An important problem in the spectral theory of metric graphs is to quantify changes in the Laplacian eigenvalues when a finite number of vertex conditions are altered. [^1] For a survey of results when the vertex conditions are of a particularly well-studied type, we refer the reader to [6, Sec. 3.1]. Abstractly, this amounts to comparing self-adjoint extensions of a symmetric operator with finite and equal defect numbers. It was recently observed [5] that the difference of eigenvalue counting functions can be efficiently bounded in terms of a certain topological index.
The relevant index, introduced by Duistermaat in [10], takes three Lagrangian subspaces as its input. In the metric graph setting these correspond to the two sets of vertex conditions under consideration and also the Dirichlet conditions at the affected vertices. Duistermaat’s definition is somewhat indirect, as it involves an arbitrary choice of a fourth Lagrangian plane that is transversal to the three planes of interest. The proof that the resulting index is independent of this choice relies on the Maslov index for paths of Lagrangian subspaces, see [10, Eq. (2.16)]. The Maslov index is an invariant depending on the homotopy class of the path (not just its endpoints), and hence it is inherently more complicated than an index depending on fixed Lagrangian subspaces.
The principal aim of this note is to define the Duistermaat index directly, without recourse to the Maslov index. We do this by giving a set of axioms (inspired by the approach of [9]) as well as a formula requiring nothing more than linear algebra. It is the set of axioms that turns out to be most useful in computation, allowing us to give short and simple proofs for many known and some unknown properties of the Duistermaat index.
One application of the developed theory is the following curious identity and its generalizations: for any invertible $n\times n$ Hermitian matrices $A$ and $B$,
$n_{-}\big(A-B\big)-n_{-}\big(B^{-1}-A^{-1}\big)=n_{-}(A)-n_{-}(B),$ (1.1)
where $n_{-}(X)$ denotes the number of negative eigenvalues (also known as the Morse index) of the matrix $X$. It is natural to wonder what happens if either $A$ or $B$ fails to be invertible. Some special cases are described Proposition 5.1, when $\ker A\subseteq\ker B$ or vice versa. The following version, however, is valid for any $n\times n$ Hermitian matrices $A$ and $B$:
$n_{-}\big(A-B\big)=n_{-}(A)-n_{-}(B)+\operatorname{\iota}({\mathcal{G}_{0}},\mathcal{G}_{B},\mathcal{G}_{A}),$ (1.2)
where $\mathcal{G}_{M}$ stands for the graph of the matrix $M$,
$\mathcal{G}_{M}:=\big\{(x,Mx):x\in\mathbb{C}^{n}\big\},$ (1.3)
and $\operatorname{\iota}$ is the Duistermaat index, which we are about to define.
@@
An important example of a Lagrangian plane is the graph of a Hermitian matrix, as in (1.3). In particular, the “horizontal” plane,
@@
The Duistermaat index [10, 24] is an integer symplectic invariant of a triple of Lagrangian planes, describing their relative position. Intuitively, it can be thought of as the maximal dimension of a subspace of $\mathcal{L}_{3}$ that lies between $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$. The formal definition we propose differs from that of [10] by taking an axiomatic approach.
@@
There exists a unique function $\operatorname{\iota}:\Lambda(n)^{3}\to\mathbb{Z}$, called the Duistermaat index, that satisfies the following properties.
@@
    $\operatorname{\iota}({\mathcal{G}_{0}},\mathcal{G}_{A},{\mathcal{G}_{\infty}})=n_{-}(A).$ (1.8)
@@
The above index takes values between 0 and $n$:
$0\leq\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})\leq n\qquad\text{for any }\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}.$ (1.11)
@@
To motivate the terminology in (3), we recall the coboundary operator $\delta$ (see [13, Chapter 3]) which maps functions of k1k-1 arguments to functions of $k$ argument by
$\big(\delta\phi\big)(x_{1},\ldots,x_{k})=\sum_{j=1}^{k}(-1)^{k-j}\phi(x_{1},\ldots,\widehat{x}_{j},\ldots,x_{k}),$ (1.12)
where a hat indicates the argument has been omitted. Then (1.10) says that the coboundary of $\operatorname{\iota}$ is zero, i.e., $\operatorname{\iota}$ is a cocycle.
@@
Theorem 1.1 will be established in Section 2 by supplying an explicit and self-contained, if somewhat cumbersome, formula for the index. However, the axioms already provide us with a computational algorithm for the index, which also establishes its uniqueness. First we reduce the computation to the case when $\mathcal{L}_{3}$ is transversal to both $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ by using the cocycle identity with some $\mathcal{L}_{4}$ that is transversal to the other three. We can then find a symplectic transformation that maps $\mathcal{L}_{3}$ to $\mathcal{G}_{\infty}$} and $\mathcal{L}_{1}$ to $\mathcal{G}_{0}$}. Since this transformation preserves transversality, the transformed $\mathcal{L}_{2}$ can be represented as the graph of a Hermitian matrix, and (1.8) can be used.
To give an illustration of Theorem 1.1, we prove a useful formula that generalizes (1.8) and also proves (1.2) by setting $A=0$ and relabelling $C$ as $A$.
@@
In the symplectic space $\mathbb{C}\oplus\mathbb{C}$, consider the Lagrangian planes $\mathcal{G}_{A}$, $\mathcal{G}_{B}$, and $\mathcal{G}_{C}$, where $A,B,C\in\mathbb{R}$ and $\mathcal{G}_{H}:=\{(z,Hz)\colon z\in\mathbb{C}\}$. Theorem 1.4 immediately yields
@@
These results (with distinct $A$, $B$ and $C$) are illustrated in Figure 1.
@@
The plan for the rest of the note is as follows: in Section 2 we show that an index satisfying the conditions in Theorem 1.1 exists and in Section 3 we establish some further properties of this index. In Section 4 we interpret the Duistermaat index as the minimal value of the Maslov index [^2] Our definition of the Maslov index, which is reviewed in Section 4, follows the normalization convention of [24]; see [5, Sec. A.4] for a comparison with other conventions in the literature. over non-decreasing paths, namely
$\operatorname{\iota}(\mathcal{L}_{0},\mathcal{L}_{1},\mathcal{M})=\min\left\{\operatorname{Mas}_{[0,1]}\!\big(\mathcal{L}(t),\mathcal{M}\big)\Big|\ \begin{matrix}\mathcal{L}(0)=\mathcal{L}_{0},\mathcal{L}(1)=\mathcal{L}_{1},\\ t\mapsto\mathcal{L}(t)\text{\ is non-decreasing}\end{matrix}\right\}.$ (1.15)
As a corollary, see Theorem 4.8, we obtain a short proof of the Zhou–Wu–Zhu identity [24, Thm. 1.1], which expresses the so-called Hörmander index as a difference of Duistermaat indices. Finally, in Section 5, we prove Theorem 1.4, which we use to deduce (1.1) and its extensions from equation (1.2). We also give an elementary derivation of (1.1) using the Haynsworth inertia formula.
@@ Acknowledgements
This material is based upon work (G.B.) supported by a grant from the Institute for Advanced Study (IAS), School of Mathematics. G.B. also acknowledges support from the NSF Grants DMS-2247473, DMS–2510345 and from the Association of Former Students of Texas A&M University (through the Faculty Development Leave program that supported the leave to IAS). G.C. acknowledges the support of NSERC grant RGPIN-2017-04259. Y.L. was supported by the NSF grant DMS-2106157, and would like to thank the Courant Institute of Mathematical Sciences at NYU and especially Prof. Lai-Sang Young for their hospitality. S.S. was supported in part by the NSF Grants DMS-2418900 and DMS–2510344, Simons Foundation grant MP–TSM–00002897, grant no. $2024154$ from the U.S.-Israel Binational Science Foundation Jerusalem, Israel, and by the Office of the Vice President for Research & Economic Development (OVPRED) at Auburn University through the Research Support Program grant.
2. Existence of the Duistermaat index
In Remark 1.3 we observed that there is at most one function $\operatorname{\iota}$ satisfying the conditions of Theorem 1.1. The following result establishes existence of such an $\operatorname{\iota}$.
@@ 2. Existence of the Duistermaat index
$\Omega(x_{1},x_{2},x_{3};y_{1},y_{2},y_{3})$ $=\omega(x_{1},y_{2})+\omega(x_{2},y_{3})+\omega(x_{3},y_{1})$ (2.1)
$\qquad-\omega(x_{2},y_{1})-\omega(x_{3},y_{2})-\omega(x_{1},y_{3}).$
@@
This is not the same as the definition of $\operatorname{\iota}$ given in [10, 24], which requires choosing an arbitrary fourth Lagrangian plane. However, since that definition satisfies the axioms given in Theorem 1.1, it must agree with the index defined by (2.2). One benefit of equation (2.2) is that $\operatorname{\iota}$ is evidently a symplectic invariant. This definition is not optimal for practical computations, however, since $\Omega$ is defined on an $3n$-dimensional space. One also notices a lack of symmetry caused by the $\dim\mathcal{L}_{1}\cap\mathcal{L}_{3}$ term. Both of these concerns are addressed by a more practical (and more symmetric) formula given in Proposition 2.3 below.
@@
$Z={\begin{pmatrix}X\\ Y\end{pmatrix}}\colon\mathbb{C}^{n}\to\mathbb{C}^{n}\oplus\mathbb{C}^{n}$ (2.3)
whose range is $\mathcal{M}$. Moreover, $\mathcal{M}$ is Lagrangian if and only if $X^{*}Y=Y^{*}X$.
@@
are independent of the choice of frame. The set $E(\mathcal{M})$ is finite, since $\det(X+\epsilon Y)$ is a polynomial in $\epsilon$ that is not identically zero. [^3] Setting $\epsilon=i$ we have $\ker(X+iY)=\ker(X+iY)^{*}(X+iY)=\ker(X^{*}X+Y^{*}Y)=\ker X\cap\ker Y=\{0\}$ because $Z$ is injective. When $\mathcal{M}$ is Lagrangian, the condition $X^{*}Y=Y^{*}X$ implies that
@@
Let $\mathcal{L}_{1}$, $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$ be Lagrangian planes and $R^{\epsilon}_{1}$, $R^{\epsilon}_{2}$ and $R^{\epsilon}_{3}$ the corresponding $\epsilon$-Robin maps, well-defined for any real $\epsilon$ outside the finite set $E_{123}:=E(\mathcal{L}_{1})\cup E(\mathcal{L}_{2})\cup E(\mathcal{L}_{3})$. Then, for the index $\operatorname{\iota}$ defined in (4.3) and any $\epsilon\in\mathbb{R}\backslash E_{123}$, we have
@@
where $\widetilde{n}_{-}(X,Y):=n_{-}(Y-X)$.
The proof of Theorem 1.1, which we give below, is an easily consequence of this proposition.
@@
We follow the proof of a similar identity in [20, Lem. 8.3] (attributed therein to Brian Feldstein). We first compute the form $\Omega$ defined in (2.1) as a Hermitian form on $\mathbb{C}^{n}\oplus\mathbb{C}^{n}\oplus\mathbb{C}^{n}$, where each $\mathbb{C}^{n}$ parametrizes the corresponding Lagrangian plane $\mathcal{L}_{j}$ via a frame $(X_{j},Y_{j})$. If, for instance,
$x_{1}=\begin{pmatrix}X_{1}u_{1}\\ Y_{1}u_{1}\end{pmatrix}\in\mathcal{L}_{1},\qquad y_{2}=\begin{pmatrix}X_{2}v_{2}\\ Y_{2}v_{2}\end{pmatrix}\in\mathcal{L}_{2},$ (2.9)
@@
$\omega(x_{1},y_{2})=\left<u_{1},(X_{1}^{*}Y_{2}-Y_{1}^{*}X_{2})v_{2}\right>_{\mathbb{C}^{n}}.$ (2.10)
Similar calculations for the other terms in (2.1) result in $\Omega$ being represented by the Hermitian matrix
@@
We now note that for $\epsilon\in\mathbb{R}\backslash E_{123}$, and $j,k\in\{1,2,3\}$,
@@
$T:=\begin{pmatrix}-I&\phantom{-}I&\phantom{-}I\\ \phantom{-}I&-I&\phantom{-}I\\ \phantom{-}I&\phantom{-}I&-I\end{pmatrix},\qquad D:=\begin{pmatrix}X_{1}+\epsilon Y_{1}&0&0\\ 0&X_{2}+\epsilon Y_{2}&0\\ 0&0&X_{3}+\epsilon Y_{3}\end{pmatrix}$, (2.14)
@@
We next observe that, by (2), (2.10), and the definition of a Lagrangian plane in (1.5),
$\ker\left(R_{k}^{\epsilon}-R_{j}^{\epsilon}\right)$ $=\ker\big(X_{j}^{*}Y_{k}-Y_{j}^{*}X_{k}\big)$
$=\big\{v\in\mathbb{C}^{n}\colon 0=\left<u,(X_{j}^{*}Y_{k}-Y_{j}^{*}X_{k})v\right>_{\mathbb{C}^{n}}\text{ for all }u\in\mathbb{C}^{n}\big\}$
$\simeq\big\{y\in\mathcal{L}_{k}\colon\omega(x,y)=0\text{ for all }x\in\mathcal{L}_{j}\big\}$
@@
therefore
$n_{0}\left(R_{k}^{\epsilon}-R_{j}^{\epsilon}\right)=\dim\mathcal{L}_{j}\cap\mathcal{L}_{k},$ (2.16)
@@
$n_{-}(\Omega)=n-\dim\mathcal{L}_{1}\cap\mathcal{L}_{3}+n_{-}\left(R_{3}^{\epsilon}-R_{2}^{\epsilon}\right)-n_{-}\left(R_{3}^{\epsilon}-R_{1}^{\epsilon}\right)+n_{-}\left(R_{2}^{\epsilon}-R_{1}^{\epsilon}\right),$ (2.18)
@@
To check the normalization condition (1.8), we apply (2.7) with $\mathcal{L}_{1}={\mathcal{G}_{0}}$, $\mathcal{L}_{2}=\mathcal{G}_{A}$ and $\mathcal{L}_{3}={\mathcal{G}_{\infty}}$. For small positive $\epsilon$ we have $R_{1}^{\epsilon}=0$, $R_{2}^{\epsilon}=A+O(\epsilon)$ and $R_{3}^{\epsilon}=\epsilon^{-1}I$, therefore
@@
Symplectic invariance of $\operatorname{\iota}$ follows from the symplectic invariance of every term in (2.2), and the cocycle property $\delta\!\operatorname{\iota}=0$ follows from (2.8) by the well-known identity $\delta^{2}\equiv 0$ for the coboundary operator, cf. Remark 1.2.

### v3

@@ 1. Introduction
The relative positions of eigenvalues of two self-adjoint operators that differ by a finite-rank perturbation in their boundary conditions is a question arising in many applications. These include the TE/TM mode analysis in graph-like photonic structures [40], stability analysis of brake orbits in Lagrangian systems [38, 37], analysis of nodal and Robin domains [13, 3, 7] and spectral minimal partitions [6, 30] on quantum graphs. From an abstract perspective, tracking eigenvalue changes due to variations in boundary conditions is equivalent to comparing self-adjoint extensions of a symmetric operator with finite and equal defect numbers. Such spectral comparison problems for differential operators may be approached via topological indices [5, 18, 22, 24]. In particular, the Maslov index [4, 39, 19, 23, 42, 47, 48] has played a major role in these investigations, and is an active area of research; see [9] for a review and [1, 7, 8, 10, 11, 15, 20, 21, 27, 32, 33, 34, 35, 36, 41] for a partial list of papers specifically focused on the finite defect setting. The main theme in these papers is to equate the Maslov index to the eigenvalue counting function of some linear differential operator, which is usually obtained by linearizing a dynamical system at an equilibrium solution. However, the Maslov index is difficult to compute in practice. It depends on a path of Lagrangian subspaces, and so without complete knowledge of this path (which amounts to having an explicit solution for the linearized system) it cannot be computed effectively.
The fundamental observation of [38, 15, 27] is that the difference of two eigenvalue counting functions can be evaluated in terms of a much simpler index that only depends on the endpoints of the path. The so-called Duistermaat index, introduced in [24], takes three Lagrangian subspaces as its input. In the metric graph setting these correspond to the two sets of vertex conditions under consideration and also the Dirichlet conditions at the affected vertices. Duistermaat’s definition is somewhat indirect, however, as it involves an arbitrary choice of a fourth Lagrangian plane that is transversal to the three planes of interest. The proof that the resulting index is independent of this choice relies on the Maslov index for paths of Lagrangian subspaces, see [24, Eq. (2.16)]. As noted above, the Maslov index is inherently more complicated than an index depending on fixed Lagrangian subspaces.
The principal aim of this note is to define the Duistermaat index directly, without recourse to the Maslov index. We do this by giving a set of axioms (inspired by the approach of [19]) as well as a formula requiring nothing more than linear algebra. It is the set of axioms that turns out to be most useful in practice, allowing us to give short and simple proofs for many known and some unknown properties of the Duistermaat index.
@@
An important example of a Lagrangian plane is the graph of a Hermitian matrix $M$,
$\mathcal{G}_{M}:=\big\{(x,Mx):x\in\mathbb{C}^{n}\big\}.$ (1.3)
In particular, the “horizontal” plane,
@@
The Duistermaat index is an integer symplectic invariant of a triple of Lagrangian planes, describing their relative position. Intuitively, it can be thought of as the maximal dimension of a subspace of $\mathcal{L}_{3}$ that lies between $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$. The formal definition we propose differs from that of [24] by taking an axiomatic approach.
@@
There exists a unique function $\operatorname{\iota}:\Lambda(n)^{3}\to\{0,1,\ldots,n\}$, called the Duistermaat index, that satisfies the following properties.
@@
    $\operatorname{\iota}({\mathcal{G}_{0}},\mathcal{G}_{A},{\mathcal{G}_{\infty}})=n_{-}(A),$ (1.6)
    where $n_{-}(A)$ denotes the number of negative eigenvalues (also known as the Morse index) of the matrix $A$.
@@
The Duistermaat index as defined in [24] and [51] satisfies the axioms in Theorem 1.1 and therefore coincides with our version.
@@
To motivate the terminology in (3), we recall the coboundary operator $\delta$ (see [28, Chapter 3]) which maps functions of k1k-1 arguments to functions of $k$ arguments by
$\big(\delta\phi\big)(\ell_{1},\ldots,\ell_{k})=\sum_{j=1}^{k}(-1)^{k-j}\phi(\ell_{1},\ldots,\widehat{\ell}_{j},\ldots,\ell_{k}),$ (1.9)
where a hat indicates the argument has been omitted. Then (1.8) says that the coboundary of $\operatorname{\iota}$ is zero, i.e., $\operatorname{\iota}$ is a cocycle. $\Diamond$
Theorem 1.1 will be established in Section 2 by supplying an explicit and self-contained, if somewhat cumbersome, formula for the index. A more practical formula is the following extension of (1.6).
@@
One application of (1.10) is the following curious identity: for any invertible $n\times n$ Hermitian matrices $A$ and $B$,
$n_{-}\big(A-B\big)-n_{-}\big(B^{-1}-A^{-1}\big)=n_{-}(A)-n_{-}(B).$ (1.11)
Non-invertible $A$ or $B$ are discussed in Section 5. We also point out that (1.10) can be used for arbitrary $\mathcal{L}_{1}$, $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$ using symplectic invariance (1.7) and a symplectic transformation $\mathsf{S}$ which makes all $\mathsf{S}\mathcal{L}_{j}$ transversal to $\mathcal{G}_{\infty}$, for example
$\mathsf{S}=\begin{pmatrix}I&\epsilon I\\ 0&I\end{pmatrix}$ (1.12)
with generic $\epsilon$, see the discussion following (2.4).
@@
In the symplectic space $\mathbb{C}\oplus\mathbb{C}$, consider three Lagrangian plane, $\mathcal{G}_{A}:=\{(z,Az)\colon z\in\mathbb{C}\}$ and the similarly defined $\mathcal{G}_{B}$ and $\mathcal{G}_{C}$, with the “slopes” $A,B,C\in\mathbb{R}$. Theorem 1.4 immediately yields
@@
These results (with distinct $A$, $B$ and $C$) are illustrated in Figure 1. The index $\operatorname{\iota}$ can be evaluated by rotating the line $\mathcal{G}_{A}$ counterclockwise to $\mathcal{G}_{B}$ and counting the number of times (0 or 1 in this low-dimensional case) that it intersects $\mathcal{G}_{C}$. The borderline cases are treated according to (1.13): $\mathcal{G}_{A}=\mathcal{G}_{C}$ counts as an intersection but $\mathcal{G}_{B}=\mathcal{G}_{C}$ does not; if $\mathcal{G}_{A}=\mathcal{G}_{B}$, the index is 0 for any $\mathcal{G}_{C}$. This “dynamic” perspective readily generalizes to the case when one or more of the Lagrangian planes is vertical, as well as to Lagrangian planes in $\mathbb{C}^{n}\oplus\mathbb{C}^{n}$ where the intersections are counted with their dimension. $\Diamond$
@@
The intuition gained in Example 1.5 allows us to interpret the Duistermaat index as the minimal value of the Maslov index over non-decreasing paths. This interpretation highlights the usefulness of the Duistermaat index, since computation of the Maslov index is far more difficult than applying an algebraic identity such as (1.10).
Theorem 1.6.
$\operatorname{\iota}(\mathcal{L}_{0},\mathcal{L}_{1},\mathcal{M})=\min\left\{\operatorname{Mas}_{[0,1]}\!\big(\mathcal{L}(t),\mathcal{M}\big)\Big|\ \begin{matrix}\mathcal{L}(0)=\mathcal{L}_{0},\,\mathcal{L}(1)=\mathcal{L}_{1},\\ t\mapsto\mathcal{L}(t)\text{\ is non-decreasing}\end{matrix}\right\},$ (1.14)
with the minimum attained by some (non-unique) path $\mathcal{L}(t)$. Furthermore, any path $\mathcal{L}(t)$ that attains the minimum for one $\mathcal{M}$ will attain the minimum for all $\mathcal{M}$.
Here, $\operatorname{Mas}_{[0,1]}\!\big(\mathcal{L}(t),\mathcal{M}\big)$ is the Maslov index as defined in (4.9); the term non-decreasing is defined in Section 4.2.
The plan for the rest of the note is as follows: in Section 2 we show that an index satisfying the conditions in Theorem 1.1 exists and is unique; in Section 3 we establish some further properties of this index. In Section 4 we prove Theorem 1.6, providing along the way a short proof (see Theorem 4.6) of the Zhou–Wu–Zhu identity [51, Thm. 1.1], which expresses the so-called Hörmander index as a difference of Duistermaat indices. Finally, in Section 5, we prove (1.11) and its extensions. We also give an elementary derivation of (1.11) using the Haynsworth inertia formula.
Notation
We use $x,y$ to denote vectors in $\mathbb{C}^{n}$ and $u,v$ to denote vectors in $\mathbb{C}^{n}\oplus\mathbb{C}^{n}$. To save space we write $u=(x,y)$ for the vector $u=\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)\in\mathbb{C}^{n}\oplus\mathbb{C}^{n}$ with components $x,y\in\mathbb{C}^{n}$. Similarly, we abbreviate $(X,Y)=\left(\begin{smallmatrix}X\\ Y\end{smallmatrix}\right)$ for the $(2n\times n)$-matrix with $(n\times n)$-matrix components $X,Y$. We write $\ker M=\{x\in\mathbb{C}^{n}:Mx=0\}$ for the kernel of an $(n\times n)$-matrix $M$ and $\mathcal{X}\simeq\mathcal{Y}$ when subspaces $\mathcal{X}$ and $\mathcal{Y}$ are isomorphic.
@@ Acknowledgements
This material is based upon work (G.B.) supported by a grant from the Institute for Advanced Study (IAS), School of Mathematics. G.B. also acknowledges support from the NSF Grants DMS-2247473, DMS–2510345 and from the Association of Former Students of Texas A&M University (through the Faculty Development Leave program that supported the leave to IAS). G.C. acknowledges the support of NSERC grants RGPIN-2017-04259 and RGPIN-2025-06186. Y.L. was supported by the NSF grant DMS-2106157, and would like to thank the Courant Institute of Mathematical Sciences at NYU and especially Prof. Lai-Sang Young for their hospitality. S.S. was supported in part by the NSF Grants DMS-2418900 and DMS–2510344, Simons Foundation grant MP–TSM–00002897, grant no. $2024154$ from the U.S.-Israel Binational Science Foundation Jerusalem, Israel, and by the Office of the Vice President for Research & Economic Development (OVPRED) at Auburn University through the Research Support Program grant.
The authors are extremely grateful to an anonymous referee who read the paper thoroughly and suggested numerous corrections and clarifications, substantially improving the manuscript.
2. Existence and uniqueness of the Duistermaat index
In Theorem 2.1 we define a function $\operatorname{\iota}$ that satisfies the conditions of Theorem 1.1. Uniqueness then follows from the observation that these conditions provide a prescription for the computation of $\operatorname{\iota}$; see Section 2.2 for details.
2.1. Existence of the index
@@
$\Omega(u_{1},u_{2},u_{3};v_{1},v_{2},v_{3})$ $=\omega(u_{1},v_{2})+\omega(u_{2},v_{3})+\omega(u_{3},v_{1})$ (2.1)
$\qquad-\omega(u_{2},v_{1})-\omega(u_{3},v_{2})-\omega(u_{1},v_{3}).$
@@
This is more direct than the definition of $\operatorname{\iota}$ given in [24, 51], which requires choosing an arbitrary fourth Lagrangian plane. One benefit of equation (2.2) is that $\operatorname{\iota}$ is evidently a symplectic invariant. This definition is not optimal for practical computations, however, since $\Omega$ is defined on a $3n$-dimensional space. One also notices a lack of symmetry caused by the $\dim\mathcal{L}_{1}\cap\mathcal{L}_{3}$ term. Both of these concerns are addressed by a more practical (and more symmetric) formula (2.7) for the index given in Proposition 2.3 below. $\Diamond$
@@
$Z=\big(X,Y\big)\colon\mathbb{C}^{n}\to\mathbb{C}^{n}\oplus\mathbb{C}^{n}$ (2.3)
whose range is $\mathcal{M}=\{(Xx,Yx):x\in\mathbb{C}^{n}\}$. Moreover, $\mathcal{M}$ is Lagrangian if and only if $X^{*}Y=Y^{*}X$, as easily seen from (1.1) and (1.2).
@@
are independent of the choice of frame. When $\mathcal{M}$ is Lagrangian, the set $E(\mathcal{M})$ is finite, since $\det(X+\epsilon Y)$ is a polynomial in $\epsilon$ that is not identically zero. [^1] Setting $\epsilon=i$, we have $\ker(X+iY)=\ker(X+iY)^{*}(X+iY)=\ker(X^{*}X+Y^{*}Y)=\ker X\cap\ker Y=\{0\}$, using $X^{*}Y=Y^{*}X$ and the injectivity of $Z$, so we conclude that $\det(X+iY)\neq 0$. The Lagrangian condition $X^{*}Y=Y^{*}X$ also implies that
@@
Let $\mathcal{L}_{1}$, $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$ be Lagrangian planes and $R^{\epsilon}_{1}$, $R^{\epsilon}_{2}$ and $R^{\epsilon}_{3}$ the corresponding $\epsilon$-Robin maps, for any real $\epsilon$ outside the finite set $E_{123}:=E(\mathcal{L}_{1})\cup E(\mathcal{L}_{2})\cup E(\mathcal{L}_{3})$. Then, for the index $\operatorname{\iota}$ defined in (2.2) and any $\epsilon\in\mathbb{R}\backslash E_{123}$, we have
@@
where $\widetilde{n}_{-}(X,Y):=n_{-}(Y-X)$. In particular, the right-hand side of (2.7), when defined, is independent of $\epsilon$.
@@
We follow the proof of a similar identity in [46, Lem. 8.3] (attributed therein to Brian Feldstein). We first compute the form $\Omega$ defined in (2.1) as a Hermitian form on $\mathbb{C}^{n}\oplus\mathbb{C}^{n}\oplus\mathbb{C}^{n}$, where each $\mathbb{C}^{n}$ parametrizes the corresponding Lagrangian plane $\mathcal{L}_{j}$ via a frame $(X_{j},Y_{j})$. If, for instance,
$u=(X_{1}x,Y_{1}x)\in\mathcal{L}_{1},\qquad v=(X_{2}y,Y_{2}y)\in\mathcal{L}_{2},$ (2.9)
@@
$\omega(u,v)=\big\langle x,(X_{1}^{*}Y_{2}-Y_{1}^{*}X_{2})y\big\rangle_{\mathbb{C}^{n}}.$ (2.10)
Similar calculations for the other terms in (2.1) result in $\Omega$ being represented by the formula
$\Omega(u_{1},u_{2},u_{3};v_{1},v_{2},v_{3})=\big\langle(x_{1},x_{2},x_{3}),W(y_{1},y_{2},y_{3})\big\rangle_{\mathbb{C}^{n}\oplus\mathbb{C}^{n}\oplus\mathbb{C}^{n}},$
where $u_{j}=(X_{j}x_{j},Y_{j}x_{j}),v_{j}=(X_{j}y_{j},Y_{j}y_{j})$, $j\in\{1,2,3\}$, with the Hermitian matrix
@@
We now note that for $\epsilon\in\mathbb{R}\backslash E_{123}$ and $j,k\in\{1,2,3\}$,
@@
$T:=\begin{pmatrix}-I&\phantom{-}I&\phantom{-}I\\ \phantom{-}I&-I&\phantom{-}I\\ \phantom{-}I&\phantom{-}I&-I\end{pmatrix},\qquad D:=\begin{pmatrix}X_{1}+\epsilon Y_{1}&0&0\\ 0&X_{2}+\epsilon Y_{2}&0\\ 0&0&X_{3}+\epsilon Y_{3}\end{pmatrix}$ (2.14)
@@
We next observe that, by (2.1) and invertibility of $X_{k}+\epsilon Y_{k}$,
$\ker\left(R_{k}^{\epsilon}-R_{j}^{\epsilon}\right)\simeq\ker\big(X_{j}^{*}Y_{k}-Y_{j}^{*}X_{k}\big).$ (2.15)
We then have
$\ker\big(X_{j}^{*}Y_{k}-Y_{j}^{*}X_{k}\big)$ $=\big\{y\in\mathbb{C}^{n}\colon 0=\left<x,(X_{j}^{*}Y_{k}-Y_{j}^{*}X_{k})y\right>_{\mathbb{C}^{n}}\text{ for all }x\in\mathbb{C}^{n}\big\}$
$\simeq\big\{v\in\mathcal{L}_{k}\colon\omega(u,v)=0\text{ for all }u\in\mathcal{L}_{j}\big\}$
@@
where we used (2.10) to get to the second line and the definition of a Lagrangian plane in (1.2) to get to the third. We can now conclude that
$n_{0}\big(R_{k}^{\epsilon}-R_{j}^{\epsilon}\big)=\dim\mathcal{L}_{j}\cap\mathcal{L}_{k},$ (2.17)
@@
$n_{-}(\Omega)=n-\dim\mathcal{L}_{1}\cap\mathcal{L}_{3}+n_{-}(R_{3}^{\epsilon}-R_{2}^{\epsilon})-n_{-}(R_{3}^{\epsilon}-R_{1}^{\epsilon})+n_{-}(R_{2}^{\epsilon}-R_{1}^{\epsilon}),$ (2.19)
@@
To check the normalization condition (1.6), we apply (2.7) with $\mathcal{L}_{1}={\mathcal{G}_{0}}$, $\mathcal{L}_{2}=\mathcal{G}_{A}$ and $\mathcal{L}_{3}={\mathcal{G}_{\infty}}$. For small positive $\epsilon$ we have $R_{1}^{\epsilon}=0$, $R_{2}^{\epsilon}=A(I+\epsilon A)^{-1}=A+o(1)$ and $R_{3}^{\epsilon}=\epsilon^{-1}I$, therefore $R_{3}^{\epsilon}-R_{2}^{\epsilon}=\epsilon^{-1}(I-\epsilon A+o(\epsilon))$ and so
@@
Symplectic invariance of $\operatorname{\iota}$ follows from the symplectic invariance of every term in (2.1), and the cocycle property $\delta\!\operatorname{\iota}=0$ follows from (2.8) by the well-known identity $\delta^{2}\equiv 0$ for the coboundary operator, which is easily verified from (1.9).